OpenStudy (anonymous):

Can someone show me the steps to: (a^2+5a+4 / a^3) divided by (a^2 + 3a + 2 / a^2 - 2a)

OpenStudy (mertsj):

\[\frac{a^2+5a+4}{a^3}\div\frac{a^2+3a+2}{a^2-2a}\]

OpenStudy (mertsj):

\[\frac{(a+4)(a+1)}{a^3}\times\frac{a(a-2)}{(a+2)(a+1)}\]

OpenStudy (mertsj):

Cancel the common factors

OpenStudy (anonymous):

\[\frac{ (a +4) }{ (a^3) } \div \frac{ a(a-2) }{ (a+12) }\]

OpenStudy (anonymous):

oops, that's suppose to be a multiplication sign, not division.